# Implicit Gaussian process representation of vector fields over arbitrary latent manifolds

### v1

@@ A.5 Binary Classification
To evaluate whether the RVGP reconstruction contained relevant information about cognitive function, we tested whether healthy controls and Alzheimer’s disease patients could be accurately classified from the ratio of source to sinks and the ratio of clockwise to anti-clockwise spirals. We tested the classification accuracy using 3 seed brain regions and compared the accuracy from each reconstructed approach (RVGP, spline, linear interpolation) against the high-density reference EEG. For binary classification, we used a linear support vector machine and computed the classification accuracy and ROC curves for each approach separately using 10-fold cross-validation. We trained two one-vs-all classifiers to binarily categorize samples pooled from all 3 brain regions and across subjects into each class (AD or healthy controls). For each fold of training, we fit the posterior distribution of the scores to 90% of the data and used this to determine the probability of each sample belonging to each class in the test set (10%). To handle class imbalance, we repeated the classification 500 times, with each iteration trained on a random subset of 100 AD and 100 healthy control samples. The AUC and ROC reported in the results were calculated from the distribution estimated over 500 iterations.
@@ A.6 Visualisation
For visualisation, the $x,y,z$ coordinates for each channel location were flattened onto a 2D grid using multi-dimensional scaling to most accurately preserve the local distances between nodes across the scalp. For plotting, we use the 2D tangent vectors of the phase field and streamlines computed using the streamslice function in MATLAB.
@@ A.7 Data Collection
Data was collected in the UKDRI Care Research & Technology Centre at the Micheal Uren Hub in London from 2022-2023. AD exclusion criteria were a previous history of head injury or any other neurological condition. All patients had a clinical diagnosis of AD, with one subject retrospectively excluded after a change in diagnosis to frontal-temporal lobe dementia (FTLD). Subjects on AD medication were asked to abstain on the day of testing and to avoid caffeine. 11 subjects were excluded from EEG analysis due to low signal quality leaving 61 subjects total (33 AD, 28 controls).

### v2

@@ A.5 Binary Classification
To evaluate whether the RVGP reconstruction contained relevant information about cognitive function, we tested whether healthy controls and Alzheimer’s disease patients could be accurately classified by the ratio of source to sinks and the ratio of clockwise to anti-clockwise spirals. We tested the classification accuracy using three seed brain regions and compared the accuracy from each reconstructed approach (RVGP, spline, linear interpolation) against the high-density reference EEG. For binary classification, we used a linear support vector machine and computed the accuracy and ROC curves for each approach separately using 10-fold cross-validation. We trained two one-vs-all classifiers to categorise pooled samples from all 3 brain regions and across subjects into each class (AD or healthy controls). For each training epoch, we fit the posterior distribution of the scores to 90% of the data and used this to determine the probability of each sample belonging to each class in the test set (10%). To handle class imbalance, we repeated the classification 500 times, with each iteration trained on a random subset of 100 AD and 100 healthy control samples. The AUC and ROC reported in the results were calculated from the distribution estimated over 500 iterations.
@@ A.6 Visualisation
For visualisation, the $x,y,z$ coordinates for each channel location were flattened onto a 2D grid using multi-dimensional scaling to most accurately preserve the local distances between nodes across the scalp. For plotting, we use the 2D tangent vectors of the phase field and streamlines computed using the ’streamslice’ function in MATLAB.
@@ A.7 Data Collection
Data was collected in the UKDRI Care Research & Technology Centre at the Micheal Uren Hub in London from 2022-2023. AD exclusion criteria were a previous history of head injury or any other neurological condition. All patients had a clinical diagnosis of AD, with one subject retrospectively excluded after a change in diagnosis to frontal-temporal lobe dementia (FTLD). Subjects on AD medication were asked to abstain on the day of testing and to avoid caffeine. 11 subjects were excluded from EEG analysis due to low signal quality, leaving 61 subjects total (33 AD, 28 controls).
Figure S1: Reconstruction of spatiotemporal wave patterns in 32-channel human EEG. Reconstruction accuracy, measured by the preservation of singularities. The accuracy is 10-fold lower than using 64-channel EEG, meaning that predictive power is lost for all methods at this resolution.
Figure S2: Superresolution for sparse training data. Same as in Fig. 2A, but for points placed at an average distance of 5% of manifold diameter.
Figure S3: Training and prediction for off-manifold points. To test the regularity in our method, we took a distribution of on-manifold points, spaced approximately 3% of manifold diameter, and added additive Gaussian noise to push them off the manifold. We increased the noise until the two dominant dimensions of the tangent space approximations explained at least 80% of the variance, amounting to a standard deviation of approximately 5% of the manifold diameter. We repeated the experiment twice, once with noise affecting only test data points and once adding noise to all data points. We trained the model as described in the main text in Section 5.1. Parameters used: $k=50$.